For the velocity components given as : u = ay sin xy, v = ax sin xy. Obtain an expression for the velocity potential function.
1
Expert's answer
2020-12-10T09:33:52-0500

If "\overrightarrow{v}" denotes the flow velocity, then for a velocity potential function "\Phi", "\overrightarrow{v}" can be represented as the gradient of a scalar function "\Phi" :


"{\displaystyle \overrightarrow{v} =\nabla \Phi \ ={\frac {\partial \Phi }{\partial x}}\mathbf {i} +{\frac {\partial \Phi }{\partial y}}\mathbf {j} +{\frac {\partial \Phi }{\partial z}}\mathbf {k} \,.}"


In our case, "\frac {\partial \Phi }{\partial x}=u=ay\sin xy,\ \ \frac {\partial \Phi }{\partial y}=v=ax\sin xy."


Then "\Phi(x,y)=\int ay\sin xy dx=-a\cos xy+C(y)", and therefore,


"\frac {\partial \Phi }{\partial y}=ax\sin xy+C'(y)".


It follows that "ax\sin xy+C'(y)=ax\sin xy", and consequently, "C'(y)=0". Then "C(y)=C", and we conclude that


"\Phi(x,y)=-a\cos xy+C."